# Dynamical discontinuities in repeated weak measurements revealed by complex weak values

### v1

@@ V Analysis of the Discontinuity
In second-order phase transitions, such as those in the Ising model [niss2005history] or the superconducting transition [kiometzis1994critical], the behavior of observables near the critical point is characterized by critical exponents that are universal across systems belonging to the same universality class. Typically, near the critical point of a control parameter $P$, for instance, the temperature, the magnetic field, or the interaction strength in measurement-induced phase transitions [poboiko2024measurement], a characteristic physical quantity $F$ (such as the heat capacity or the magnetic susceptibility) diverges following a power law in the vicinity of the critical point:
@@
In the protocol studied in this article, the post-selected polar angle $\phi$ acts as the control parameter, governing a discontinuity in the expectation value of the Pauli matrix $\hat{\sigma}_{x}$ at the critical point $\phi_{c}$. This discontinuity emerges after multiple iterations of the protocol. The corresponding relaxation time, defined as the time required for the system to reach its steady state, diverges at $\phi_{c}$, since at the critical point the meter dynamics become effectively unitary and no single eigenstate (fixed point) dominates. Although the discontinuity observed in Fig. 2 might at first appear reminiscent of a first-order phase transition, the absence of a direct jump between two stable fixed points, and the emergence of unitary dynamics precisely at the critical point, indicate a resemblance to a second-order phase transition. The divergence of the relaxation time [dziarmaga2010dynamics], demonstrated below, signifies critical slowing down, a defining feature of second-order phase transitions. Thus, the relaxation time $\tau$ serves as a key quantity for characterizing the critical behavior and determining the associated critical exponent, as well as for probing the universality of the transition. This section demonstrates that the relaxation time associated with the proposed protocol universally follows the power-law behavior expressed in:
@@
The precise definition of the relaxation time depends on the type of dynamics. Here, the dynamics proceed in discrete time steps, so $\tau$ can be defined in terms of the spectrum of the operator driving the dynamics, specifically the Kraus operator. This analysis determines the critical exponent from the relaxation dynamics, in analogy with dynamic Markovian processes [prinz2011markov, levin2017markov], where the key quantity is the ratio between the two leading eigenvalues of the Kraus operator. To quantify the convergence rate of the iterated Kraus map toward its fixed point, the relaxation time is defined in terms of the map’s spectral properties. Consider the unnormalized meter state after $n$ applications of the Kraus operator:
@@
Denoting by $|\lambda_{1}|$ and $|\lambda_{2}|$ the eigenvalues of the Kraus operator $\hat{K}$ with the largest moduli, such that $|\lambda_{1}|>|\lambda_{2}|$, the dominant diagonal term scales as $|\lambda_{1}|^{2n}$, while the next-to-leading term scales as $|\lambda_{2}|^{2n}$.
Consequently, the ratio between subdominant and dominant contributions decays as $(|\lambda_{2}|/|\lambda_{1}|)^{2n}$.
Matching this discrete decay to an effective exponential relaxation $e^{-n/\tau}$ defines the relaxation time $\tau$:
@@
For any system evolving under the dynamics dictated by the weak measurement protocol, the eigenvalues of the Kraus operator are given by Eq. 7. We assume that the meter operator $\hat{O}_{A}$ is non-degenerate. In this case, let us denote the largest eigenvalue in modulus by $\lambda_{1}$ and the second-largest by $\lambda_{2}$, corresponding to the eigenvalues $o_{1}$ and $o_{2}$ of the operator $\hat{O}_{A}$, respectively. The squared moduli of these eigenvalues can then be expressed, to first order in $gt$, as
@@ VI Conclusion
Complex weak values are essential for the emergence of such dynamical discontinuities, whereas purely real weak values suppress them. The relaxation time diverges at the critical point and exhibits a scaling law that appears to be universal, independent of the protocol parameters, with a critical exponent of 1. This behavior is reminiscent of second order phase transitions, as has been demonstrated analytically for systems of any dimension and confirmed numerically.
A nonzero imaginary part of the weak value modifies the system’s evolution from a process that is unitary at first order in $gt$, where the Kraus operator has real eigenvalues of equal modulus at first order in $gt$, to a non-unitary dynamics featuring $N$ fixed points, with $N$ denoting the dimension of the system’s Hilbert space. Among these, one fixed point is stable, provided that the Kraus operator is non-degenerate. In this case, the dynamics drives the system toward the stable fixed point, which corresponds to the eigenvector associated with the largest eigenvalue of the Kraus operator, unless the meter is initially prepared in exactly one of the other eigenvectors corresponding to unstable fixed points. This eigenvector coincides with the largest eigenvalue of the meter operator if the imaginary part of the weak value is positive, and with the smallest eigenvalue if the imaginary part of the weak value is negative.

### v2

@@ V Analysis of the Discontinuity
In second-order phase transitions, such as those in the Ising model [37] or the superconducting transition [28], the behavior of observables near the critical point is characterized by critical exponents that are universal across systems belonging to the same universality class. Typically, near the critical point of a control parameter $P$, for instance, the temperature, the magnetic field, or the interaction strength in measurement-induced phase transitions [39], a characteristic physical quantity $F$ (such as the heat capacity or the magnetic susceptibility) diverges following a power law in the vicinity of the critical point:
@@
In the protocol studied in this article, the post-selected polar angle $\phi$ acts as the control parameter, governing a discontinuity in the expectation value of the Pauli matrix $\hat{\sigma}_{x}$ at the critical point $\phi_{c}$. This discontinuity emerges after multiple iterations of the protocol. The corresponding relaxation time, defined as the time required for the system to reach its steady state, diverges at $\phi_{c}$, since at the critical point the meter dynamics become effectively unitary and no single eigenstate (fixed point) dominates. Although the discontinuity observed in Fig. 2 might at first appear reminiscent of a first-order phase transition, the absence of a direct jump between two stable fixed points, and the emergence of unitary dynamics precisely at the critical point, indicate a resemblance to a second-order phase transition. The divergence of the relaxation time [12], demonstrated below, signifies critical slowing down, a defining feature of second-order phase transitions. Thus, the relaxation time $\tau$ serves as a key quantity for characterizing the critical behavior and determining the associated critical exponent, as well as for probing the universality of the transition. This section demonstrates that the relaxation time associated with the proposed protocol universally follows the power-law behavior expressed in:
@@
The precise definition of the relaxation time depends on the type of dynamics. Here, the dynamics proceed in discrete time steps, so $\tau$ can be defined in terms of the spectrum of the operator driving the dynamics, specifically the Kraus operator. This analysis determines the critical exponent from the relaxation dynamics, in analogy with dynamic Markovian processes [40, 32], where the key quantity is the ratio between the two leading eigenvalues of the Kraus operator. To quantify the convergence rate of the iterated Kraus map toward its fixed point, the relaxation time is defined in terms of the map’s spectral properties. Consider the unnormalized meter state after $n$ applications of the Kraus operator:
@@
Denoting by $|\lambda_{1}|$ and $|\lambda_{2}|$ the eigenvalues of the Kraus operator $\hat{K}$ with the largest moduli, such that $|\lambda_{1}|>|\lambda_{2}|$, the dominant diagonal term scales as $|\lambda_{1}|^{2n}$, while the next-to-leading term scales as $|\lambda_{2}|^{2n}$. Consequently, the ratio between subdominant and dominant contributions decays as $(|\lambda_{2}|/|\lambda_{1}|)^{2n}$. Matching this discrete decay to an effective exponential relaxation $e^{-n/\tau}$ defines the relaxation time $\tau$:
@@
For any system evolving under the dynamics governed by the weak measurement protocol, the eigenvalues of the Kraus operator are given by Eq. 8. We assume that the meter operator $\hat{O}_{A}$ is non-degenerate. In this case, let us denote the largest eigenvalue in modulus by $\lambda_{1}$ and the second-largest by $\lambda_{2}$, corresponding to the eigenvalues $o_{1}$ and $o_{2}$ of the operator $\hat{O}_{A}$, respectively. The squared moduli of these eigenvalues can then be expressed, to first order in $gt$, as
@@ VI Conclusion
Complex weak values are essential for the emergence of dynamical discontinuities, whereas purely real weak values suppress them. The relaxation time diverges at the critical point and exhibits a scaling law that appears to be universal, independent of the protocol parameters, with a critical exponent of 1. This behavior is reminiscent of second order phase transitions, as has been demonstrated analytically for systems of any dimension and confirmed numerically.
A nonzero imaginary part of the weak value modifies the system’s evolution from a process that is unitary at first order in $gt$, where the Kraus operator has real eigenvalues of equal modulus at first order in $gt$, to a non-unitary dynamics featuring $N$ fixed points, with $N$ denoting the dimension of the system’s Hilbert space. Among these, one fixed point is stable, provided that the Kraus operator is non-degenerate. In this case, the dynamics drives the system toward the stable fixed point, which corresponds to the eigenvector associated with the largest eigenvalue of the Kraus operator, unless the meter is initially prepared in exactly one of the other eigenvectors corresponding to unstable fixed points. This eigenvector coincides with the smallest eigenvalue of the meter operator if the imaginary part of the weak value is positive, and with the largest eigenvalue if the imaginary part of the weak value is negative.
@@
The results demonstrate that the weak value can act as a tunable parameter capable of modifying the stability structure of measurement-induced quantum dynamics. By continuously adjusting the post-selection angle, the protocol induces abrupt changes in long-time meter observables, enabling controlled manipulation of the meter state and measurement-based switching within finite-dimensional quantum systems, without requiring many-body interactions or nonlinear Hamiltonians. The asymptotic states correspond to fixed points of the associated Kraus map, providing a reliable route for state preparation. Repeated post-selected measurements generate effective non-unitary dynamics that act as dissipation, and tuning the weak value determines which fixed point the system relaxes to, realizing a form of measurement-induced dissipative control. Near the critical point, the sharp dependence of observables on the post-selection angle enhances sensitivity to small variations, suggesting applications in high-precision parameter estimation. This protocol therefore provides a versatile platform for exploring critical-like phenomena and dynamical bifurcations in analytically tractable, few-body systems, and could potentially be realized in current experimental setups with small, well-controlled quantum systems.